OpenStudy (lena772):

Differentiate h(t)=2cot^2(pi*t + 2)

ganeshie8 (ganeshie8):

chain rule

ganeshie8 (ganeshie8):

think cot(pit + 2) = x in ur head, then it looks like x^2 right ?

OpenStudy (lena772):

yes

ganeshie8 (ganeshie8):

wats derivative of x^2 ?

OpenStudy (lena772):

1

ganeshie8 (ganeshie8):

nope. use below formula :- \(\large \frac{d}{dx}(x^n) = nx^{n-1}\)

OpenStudy (lena772):

:/

OpenStudy (shamil98):

the derivative of x is 1 x^2 is 2x

ganeshie8 (ganeshie8):

\(\large \frac{d}{dx}(x^2) = 2x^{2-1} \)

OpenStudy (shamil98):

You multiply x by the power as ganeshie has shown and take the power down by 1

OpenStudy (lena772):

ok i understand

ganeshie8 (ganeshie8):

good, so we have this :- \(h(t)=2 \cot^2(pi*t + 2) \) *think \(x = \cot(pi*t + 2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2x \times \frac{d}{dt}(x)\)

ganeshie8 (ganeshie8):

replace x now

ganeshie8 (ganeshie8):

good, so we have this :- \(h(t)=2 \cot^2(pi*t + 2) \) *think \(x = \cot(pi*t + 2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2x \times \frac{d}{dt}(x)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times \frac{d}{dt}(\cot(pi*t+2))\)

ganeshie8 (ganeshie8):

you wid me still ? :o

OpenStudy (lena772):

yes

ganeshie8 (ganeshie8):

good, keep taking the derivatives. its a chain !

ganeshie8 (ganeshie8):

good, so we have this :- \(h(t)=2 \cot^2(pi*t + 2) \) *think \(x = \cot(pi*t + 2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2x \times \frac{d}{dt}(x)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times \color{red}{\frac{d}{dt}(\cot(pi*t+2))}\)

ganeshie8 (ganeshie8):

that red part u still need to work,

ganeshie8 (ganeshie8):

*think in ur head, pi*t + 2 = x then, it looks like cot(x) right ?

OpenStudy (lena772):

yea

ganeshie8 (ganeshie8):

wats derivative of cot(x) ?

ganeshie8 (ganeshie8):

derivative of cot is -cosec^2, use it there

OpenStudy (lena772):

csc ^2 x

ganeshie8 (ganeshie8):

good, so we have this :- \(h(t)=2 \cot^2(pi*t + 2) \) *think \(x = \cot(pi*t + 2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2x \times \frac{d}{dt}(x)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times \color{red}{\frac{d}{dt}(\cot(pi*t+2))}\) *think \(x = pi*t+2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times (-\csc^2x )\color{red}{\frac{d}{dt}(x)}\)

ganeshie8 (ganeshie8):

good, so we have this :- \(h(t)=2 \cot^2(pi*t + 2) \) *think \(x = \cot(pi*t + 2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2x \times \frac{d}{dt}(x)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times \color{red}{\frac{d}{dt}(\cot(pi*t+2))}\) *think \(x = pi*t+2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times (-\csc^2x )\color{red}{\frac{d}{dt}(x)}\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times (-\csc^2(pi*t+2) )\color{red}{\frac{d}{dt}(pi*t+2)}\)

ganeshie8 (ganeshie8):

one last round of derivative, then we're done.

ganeshie8 (ganeshie8):

wats derivative of \(pi*t + 2\) ?

OpenStudy (lena772):

idk

ganeshie8 (ganeshie8):

its just \(pi\)

ganeshie8 (ganeshie8):

good, so we have this :- \(h(t)=2 \cot^2(pi*t + 2) \) *think \(x = \cot(pi*t + 2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2x \times \frac{d}{dt}(x)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times \color{red}{\frac{d}{dt}(\cot(pi*t+2))}\) *think \(x = pi*t+2)\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times (-\csc^2x )\color{red}{\frac{d}{dt}(x)}\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times (-\csc^2(pi*t+2) )\color{red}{\frac{d}{dt}(pi*t+2)}\) \(\large \frac{d}{dt}(h(t))=2 \times 2\cot(pi*t+2) \times (-\csc^2(pi*t+2) \times pi \)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (lena772):

yes

ganeshie8 (ganeshie8):

good, remember this :- \(\large \frac{d}{d\color{red}{x}}(cx) = c\) \(\large \frac{d}{d\color{red}{t}}(ct) = c\)

ganeshie8 (ganeshie8):

we're done wid the differentiation btw.

OpenStudy (lena772):

yay thank you so much!